OpenStudy (anonymous):

The sum of a geometric series is -1023 and its first term is -3. If the last term is -768, what is the common ratio?

OpenStudy (anonymous):

use the sum formula of geometric series

OpenStudy (anonymous):

idk it

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Geometric_series#Sum

OpenStudy (anonymous):

\[\sum_{k=0}^{n-1}ar^k = a\frac{1 - r^n}{1 - r}\]

OpenStudy (anonymous):

\[\large Sum = \frac{a(r^n - 1)}{r - 1}\]

OpenStudy (hba):

WELCOME TO OPEN STUDY

OpenStudy (anonymous):

@hba right lol